# Quasirandom and quasisimple groups

### v2

@@ Abstract.
Key words and phrases:
Quasirandom groups, quasisimple groups.
2020 Mathematics Subject Classification:
Primary: 20D99, 20C99.
1. Introduction
@@
Fix $\varepsilon>0$. If $G$ is an $\varepsilon$-quasirandom group, then $|G/Sol(G)|\geq|G|^{\varepsilon}$. If $|G|$ is sufficiently large depending on $\varepsilon$, then $G/Sol(G)$ is a direct product of boundedly many finite simple groups of Lie type of bounded rank.
Conversely, suppose that $G$ is perfect, $|G/Sol(G)|\geq|G|^{\varepsilon}$, and that $G/Sol(G)$ is $\varepsilon$-quasirandom. If $|G|$ is sufficiently large depending on $\varepsilon$, then $G$ is $\frac{\varepsilon^{2}}{3}$-quasirandom.
For the second main result, we study the groups that are $\varepsilon$-quasirandom with a relatively large $\varepsilon$. It is a widespread belief that $(\mathrm{PSL}_{2}(q))_{q\geq 4}$ is the “most quasirandom” family of groups [23, 28], but not much is known. Not only this is actually the case, but we prove that the vast majority of highly quasirandom groups are quasisimple, shedding a good light on the issue.
@@ 3. $(c)$
The paper is structured as follows. In Section 2 we introduce the (exponential) quasirandom degree $\bm{\varepsilon}(G)$ of a finite group, and give basic results about quasisimple groups, modular and projective representations. In Section 3, we provide some highly quasirandom nonquasisimple groups: they are of affine type as in Theorem Bb. We prove the main theorems in Section 4. Fundamental to our discussion is the knowledge of the quasirandom degrees of the finite simple groups, so we collect these data in the Appendix.
2. Preliminaries
2.1. Minimal degree of a nontrivial representation
@@ 3. $(iii)$
for every finite perfect group $G$. A converse inequality, $\mathbf{D}_{0}(G)<\bm{\iota}(G)$, follows from the previous lemmas. Lev [16] proved that $\bm{\iota}(G)\leq|G|^{0.5}$ for every finite group $G$ whose order is not a prime (equality is achieved by groups of order $p^{2}$). From Corollaries C and D, we obtain $\mathbf{D}_{0}(G)<|G|^{0.2143}$ whenever $G$ is not quasisimple, and $\mathbf{D}_{0}(G)<|G|^{0.3464}$ for every finite group.
2.2. Quasirandom degree
@@
Let $G$ be a finite perfect group. The following statements are roughly equivalent, in the sense that if one statement holds for a small constant $\varepsilon>0$, then all others hold with constants that are a positive fraction of $\varepsilon$:
  • $G$ is $\varepsilon$-quasirandom;
  • every proper subgroup of $G$ has index at least $|G|^{\varepsilon}$;
  • for every subset $S$ of size at least $|G|^{1-\varepsilon}$ we have $S^{3}=G$;
  • every product-free subset of $G$ has size at most $|G|^{1-\varepsilon}$.
Let $N\lhd G$ be a proper normal subgroup. From an easy computation, we write
@@
(2.1) $|N|\>\leq\>\frac{\mathbf{D}_{0}(G/N)^{1/\bm{\varepsilon}(G)}}{|G/N|}.$
What makes (2.1) so useful is that in many occasions we barely know that $\bm{\varepsilon}(G)$ is large, but we have a good knowledge of $G/N$. The next lemma provides a gap between $\bm{\varepsilon}(G)$ and $\bm{\varepsilon}(G/N)$ when $N$ is abelian but noncentral, and it greatly simplifies the proof of Theorem B.
@@
Since $N$ is abelian, we have $N\subseteq\mathbf{C}_{G}(N)\neq G$. Combining Lemma 2.1, Corollary 2.3(ii), and (2.1), we write
@@
Let $A$ and $B$ be two finite groups, and suppose that $\bm{\varepsilon}(A)\leq\bm{\varepsilon}(B)$. Then
$\frac{\bm{\varepsilon}(A)}{2}\>\leq\>\bm{\varepsilon}(A\times B)\>\leq\>\frac{\bm{\varepsilon}(B)}{2}.$
@@
$\begin{split}\bm{\varepsilon}(A\times B)&=\frac{\log\left(\min\{\mathbf{D}_{0}(A),\mathbf{D}_{0}(B)\}\right)}{\log(|A||B|)}\\ &=\frac{\log\left(\min\{|A|^{\bm{\varepsilon}(A)},|B|^{\bm{\varepsilon}(B)}\}\right)}{\log|A|+\log|B|}\\ &\leq\bm{\varepsilon}(B)\cdot\frac{\min\{\log|A|,\log|B|\}}{\log|A|+\log|B|}\\ &\leq\frac{\bm{\varepsilon}(B)}{2}.\end{split}\,$
The proof of the other inequality is similar. ∎
Corollary 2.9.
Let $(G_{i})_{i=1}^{n}$ be finite groups and suppose $\bm{\varepsilon}(G_{1})\leq...\leq\bm{\varepsilon}(G_{n})$. Then $\bm{\varepsilon}(A_{1})/n\leq\bm{\varepsilon}(\prod_{i=1}^{n}A_{i})\leq\bm{\varepsilon}(A_{n})/n$.
2.3. Quasisimple groups
@@
The next two lemmas are essentially well known. We write down Lemma 2.11 below, which is a little more general, and will be useful later.
@@
Let $Z=\mathbf{Z}(G)$. We first show that $\mathbf{Z}(G/N)=Z/N$. It is clear that $Z/N\subseteq\mathbf{Z}(G/N)$. On the other hand, $\mathbf{Z}(G/N)$ is certainly contained in the projection of $\mathbf{Z}(G/Z)$ in $G/N$. By Grün’s Lemma [12] we have $\mathbf{Z}(G/Z)=1$, so $\mathbf{Z}(G/N)\subseteq Z/N$.
Now from the hypotheses we have $N_{0}/N\subseteq Z/N$, i.e. $N_{0}\subseteq Z$. ∎
We conclude this subsection with the following observation, which is useful for an inductive argument.
@@
$[h,g]N_{i}=[hN_{i},gN_{i}]\subseteq N_{i}.$
It follows that $[h,g]\in N_{1}\cap N_{2}=1$, and since $h\in H$ is arbitrary, we conclude $H\subseteq\mathbf{Z}(G)$. Finally, by Lemma 2.10, $G$ itself is quasisimple, against our initial assumption. ∎
2.4. Linear and projective representations
@@
For the proof of Theorem B, we need detailed informations about the $\frac{1}{4}$-quasirandom finite simple groups. The infinite families of $\frac{1}{4}$-quasirandom simple groups are
@@
$G_{2}(q)\hbox{ with }q=3(\mathrm{mod}\,6),\quad^{2}B_{2}(q^{2}),\quad^{2}G_{2}(q^{2})\,.$
The relevant informations can be recovered from [15, 24] for classical groups, and from [17, Sec. 4] for exceptional groups (see also the Appendix). We collect these data in Table 1.
@@
$\dfrac{1^{1}}{1_{1_{1}}}$ $\mathrm{PSL}_{2}(q)$ $q$ even $q-1$ $q-1$ $q-1$ $2$
[…]
$\dfrac{1^{1}}{1_{1_{1}}}$ $\mathrm{PSL}_{3}(q)$ $q\neq 2,4$ $q(q+1)$ $q(q+1)$ $q^{2}+q-1$ $3$
$\dfrac{1^{1}}{1_{1_{1}}}$ $\mathrm{Sp}_{4}(q)$ $q$ even $\dfrac{1}{2}q(q-1)^{2}$ $\dfrac{1}{2}q(q-1)^{2}$ $\dfrac{1}{2}q(q-1)^{2}$ $4$
[…]
$\dfrac{1^{1}}{1_{1_{1}}}$ ${}^{2}B_{2}(q^{2})$ $q=\sqrt{2^{2m+1}}$ $\dfrac{\sqrt{2}}{2}q\phi_{1}(q)\phi_{2}(q)$ $\dfrac{\sqrt{2}}{2}q\phi_{1}(q)\phi_{2}(q)$ $\dfrac{\sqrt{2}}{2}q\phi_{1}(q)\phi_{2}(q)$ $4$
$\dfrac{1^{1}}{1_{1_{1}}}$ ${}^{2}G_{2}(q^{2})$ $q=\sqrt{3^{2m+1}}$ $\phi_{12}(q)$ $\phi_{12}(q)$ $\phi_{12}(q)-\delta_{2}$ $7$
Table 1. The $\frac{1}{4}$-quasirandom simple groups of Lie type.
We denote by $\phi_{\ell}(q)$ the $\ell$-th cyclotomic polynomial in the variable $q$. The value of $\delta_{1}$ is $1$ if $r=2$, $q=3^{f}$, $f\geq 2$, or $r=3$, $3$ divides $q-1$, and it is 00 otherwise. The value of $\delta_{2}$ is $1$ if $r=2$, and 00 otherwise.
@@
3. Quasirandom groups of affine type
In this short section, we provide some highly quasirandom nonquasisimple groups. The key tool is Lemma 3.2 below. Before giving the results, we revise some notions from the representation theory of finite groups. Let $V$ be a finite vector space, and let $V^{*}$ be its dual space. We remark that $V^{*}\cong Hom(V,\mathbb{C}^{*})$. If $\phi:V\rightarrow V$ is a homomorphism of vector spaces, then $\phi^{*}:V^{*}\rightarrow V^{*}$ is defined by
$\phi^{*}(f)\>:=\>\phi\circ f.$
Let $K$ be a finite group. If $\rho:K\rightarrow\mathrm{GL}(V)$ is a representation of $K$, then the dual representation $\rho^{*}:K\rightarrow\mathrm{GL}(V^{*})$ is defined by
@@ 3. Quasirandom groups of affine type
4. Proof of the main theorems
4.1. Proof of Theorem A
@@
Fix $\varepsilon,\delta>0$, and let $G$ be a perfect group. Suppose that $G/Sol(G)$ is $\varepsilon$-quasirandom, and that $|G/Sol(G)|\geq|G|^{\delta}$. If $|G|$ is sufficiently large with respect to $\varepsilon$ and $\delta$, then $G$ is $\frac{\varepsilon\delta}{3}$-quasirandom.
@@
$\mathbf{D}_{0}(G)\>\geq\>c\,\mathbf{D}_{0}(G/Sol(G))^{1/2},$
where $c$ is an absolute constant. Looking back at [19], we check that $c=10^{-10}$ works. It follows that
[…]
@@
$\bm{\varepsilon}(G)\>\geq\>\frac{\bm{\varepsilon}(G/Sol(G))}{2}\cdot\frac{\log|G/Sol(G)|}{\log|G|}-\frac{10\log 10}{\log|G|}.$
@@
The main idea behind the other direction of Theorem A is the following. Under the hypothesis that $Sol(G)=1$, we prove that $G=F^{*}(G)$, the generalized Fitting subgroup of $G$. If this is not the case, then $G/F^{*}(G)$ has a nontrivial simple quotient $L$. We use the quasirandomness of $L$ to obtain an upper bound on $|G|$, while Bender’s theorem and other considerations provide a lower bound on $|F^{*}(G)|$. These inequalities fight each other, giving the desired contradiction. A similar argument will be used in a portion of the proof of Theorem B. Going into details, we start reporting the Schreier conjecture, which is now a theorem as a consequence of the Classification of the Finite Simple Groups.
@@
Let $N\lhd G$, and suppose that a section $H_{0}/N_{0}\cong L$ of $G$ is simple. Then $N\subseteq N_{0}$, or $NN_{0}=H_{0}$. In the first case $L$ appears as a section of $G/N$, while in the second case $L$ appears as a section of $N$.
@@
After Lemma 4.1, we have only to prove the first part of the theorem. Fix $\varepsilon>0$. From Corollary 2.3(i), it is clear that $|G/Sol(G)|\geq|G|^{\varepsilon}$. Moreover, $G/Sol(G)$ is again $\varepsilon$-quasirandom. Up to replacing $G$ with $G/Sol(G)$, we can assume that $G$ has a trivial solvable radical. Now it is well known that the generalized Fitting subgroup $F=F^{*}(G)$ coincides with the socle of $G$, and it is the direct product of nonabelian simple groups $T_{i}$’s. Let $F=\prod_{i}(T_{i})^{n_{i}}$, where $n_{i}\geq 1$ for each $i$. Moreover, as $\mathbf{C}_{G}(F)\subseteq F$ [1, (31.13)], we have $\mathbf{C}_{G}(F)=\mathbf{Z}(F)=1$. It follows that $G$ can be embedded as a subgroup of
@@
[…]
When $\mathbf{D}_{0}(L)$ is sufficiently large, this gives a contradiction to (4.1). We have proven that $G/Sol(G)=F^{*}(G/Sol(G))$, so $G/Sol(G)$ is an $\varepsilon$-quasirandom direct product of finite simple groups, say $G/Sol(G)=\prod_{i}(T_{i})^{n_{i}}$. From Remark 4.5, the $T_{i}$’s are all simple groups of Lie type. Let $N=\sum_{i}n_{i}$. From Corollary 2.9, we have $\varepsilon\leq\mathbf{D}_{0}(G/Sol(G))\leq N^{-1}$, so that $N\leq\varepsilon^{-1}$. Moreover, each block $(T_{i})^{n_{i}}$ has to be $\varepsilon$-quasirandom as a quotient of $G$. From Corollary 2.9 and Remark 4.5, we have
$\varepsilon\>\leq\>\bm{\varepsilon}((T_{i})^{n_{i}})\>=\>\frac{\bm{\varepsilon}(T_{i})}{n_{i}}\>\leq\>\frac{10}{n_{i}\,rank(T_{i})},$
so that $n_{i}\,rank(T_{i})\leq 10\varepsilon^{-1}$ for each $i$. ∎
@@
An attentive reader will have noticed that a bound of the type $\mathbf{D}_{0}(G)\geq c\log|G|$ is strong enough to give that $G/Sol(G)$ is the direct product of nonabelian finite simple groups.
It is natural to ask whether the condition $|G/Sol(G)|\geq|G|^{\varepsilon}$ is actually superflous in the second part of Theorem A, as it may be implied by the quasirandomness of the sole $G/Sol(G)$. The answer is no, as it can even happen that $|G/Sol(G)|$ is bounded: it is enough to consider the perfect deleted permutation module $(C_{n})^{4}\rtimes\mathrm{Alt}(5)$ when $n$ is large. The following example, where $|G/Sol(G)|$ is unbounded but too small, is due to S. Eberhard.
@@
It can be proven that $G_{d}$ is perfect, and its solvable radical is the subgroup where $g\in\mathbf{Z}(\mathrm{SL}_{2}(p))$ is a scalar matrix. In particular, $G_{d}/Sol(G_{d})$ is isomorphic to $\mathrm{PSL}_{2}(p)$, which is $0.34$-quasirandom for large $p$, but its cardinality is less than $p^{3}$.
4.2. Proof of Theorem B
We first prove the shallow version of Theorem B. This is a nice application of (2.1).

### v3

@@ Abstract
[…]
[classification=text]
1 Introduction
@@
Fix $\varepsilon>0$. If $G$ is an $\varepsilon$-quasirandom group, then $|G/\mathrm{Sol}(G)|\geq|G|^{\varepsilon}$. If $|G|$ is sufficiently large depending on $\varepsilon$, then $G/\mathrm{Sol}(G)$ is a direct product of boundedly many finite simple groups of Lie type of bounded rank.
Conversely, suppose that $G$ is perfect, $|G/\mathrm{Sol}(G)|\geq|G|^{\varepsilon}$, and that $G/\mathrm{Sol}(G)$ is $\varepsilon$-quasirandom. If $|G|$ is sufficiently large depending on $\varepsilon$, then $G$ is $\frac{\varepsilon^{2}}{3}$-quasirandom.
For the second main result, we study the groups that are $\varepsilon$-quasirandom with a relatively large $\varepsilon$. It is a widespread belief that $(\mathrm{PSL}_{2}(q))_{q\geq 4}$ is the “most quasirandom” family of groups [23, 28], but not much is known. Not only is this actually the case, but we prove that the vast majority of highly quasirandom groups are quasisimple, shedding a good light on the issue.
@@ 3. $(c)$
The paper is organized as follows. In Section 2 we introduce the (exponential) quasirandom degree $\bm{\varepsilon}(G)$ of a finite group, and give basic results about quasisimple groups, modular and projective representations. In Section 3, we provide some highly quasirandom nonquasisimple groups: they are of affine type as in Theorem B a-b. We prove the main theorems in Section 4. Fundamental to our discussion is the knowledge of the quasirandom degrees of the finite simple groups, so we collect these data in the Appendix A.
2 Preliminaries
2.1 Minimal degree of a nontrivial representation
@@
Observe that the veracity of Corollaries C and D implies that $\mathbf{D}_{0}(G)<|G|^{0.3464}$ for every finite group $G$, and $\mathbf{D}_{0}(G)<|G|^{0.2143}$ if $G$ is not quasisimple.
@@ 3. $(iii)$
for every finite perfect group $G$. A converse inequality, $\mathbf{D}_{0}(G)<\bm{\iota}(G)$, follows from the previous lemmas. Lev [16] proved that $\bm{\iota}(G)\leq|G|^{0.5}$ for every finite group $G$ whose order is not a prime (equality is achieved by groups of order $p^{2}$).
2.2 Quasirandom degree
@@
Let $G$ be a finite perfect group. The following statements are roughly equivalent, in the sense that if one statement holds for a constant $\varepsilon_{i}>0$, then all others hold with constant $\varepsilon_{j}=c_{i,j}\varepsilon_{i}$, for some $c_{i,j}>0$.
  • $G$ is $\varepsilon_{1}$-quasirandom;
  • every proper subgroup of $G$ has index at least $|G|^{\varepsilon_{2}}$;
  • for every subset $S$ of size at least $|G|^{1-\varepsilon_{3}}$ we have $S^{3}=G$;
  • every product-free subset of $G$ has size at most $|G|^{1-\varepsilon_{4}}$.
Let $G$ be a finite group and $N\lhd G$ a proper normal subgroup. We can write
@@
$|N|\>\leq\>\frac{\mathbf{D}_{0}(G/N)^{1/\bm{\varepsilon}(G)}}{|G/N|}.$ (1)
What makes (1) so useful is that in many occasions we barely know that $\bm{\varepsilon}(G)$ is large, but we have a good knowledge of $G/N$. The next lemma provides a gap between $\bm{\varepsilon}(G)$ and $\bm{\varepsilon}(G/N)$ when $N$ is abelian but noncentral, and it greatly simplifies the proof of Theorem B.
@@
Since $N$ is abelian, we have $N\subseteq\mathbf{C}_{G}(N)\neq G$. Combining Lemma 2.1, Corollary 2.3(ii), and (1), we write
@@
The following holds:
  • (i)
    If $\bm{\varepsilon}(A)\leq\bm{\varepsilon}(B)$, then $\bm{\varepsilon}(A\times B)\leq\frac{\bm{\varepsilon}(B)}{2}$;
  • (ii)
    if $(G_{i})_{i=1}^{n}$ are finite groups, then $\bm{\varepsilon}\left(\prod_{i=1}^{n}G_{i}\right)<\frac{1}{n}$.
@@
$\begin{split}\bm{\varepsilon}(A\times B)&=\frac{\log\left(\min\{\mathbf{D}_{0}(A),\mathbf{D}_{0}(B)\}\right)}{\log(|A||B|)}\\ &=\frac{\log\left(\min\{|A|^{\bm{\varepsilon}(A)},|B|^{\bm{\varepsilon}(B)}\}\right)}{\log|A|+\log|B|}\\ &\leq\bm{\varepsilon}(B)\cdot\frac{\min\{\log|A|,\log|B|\}}{\log|A|+\log|B|}.\end{split}\,$
Point (i) follows. For (ii), observe that there is a group, say $G_{n}$, satisfying $\log|G_{n}|\leq\frac{1}{n}\sum_{i=1}^{n}\log|G_{i}|$. It follows from the previous computation that
$\bm{\varepsilon}\left(\prod_{i=1}^{n}G_{i}\right)<1\cdot\frac{\min\{\sum_{i=1}^{n-1}\log|G_{i}|,\log|G_{n}|\}}{\sum_{i=1}^{n}\log|G_{n}|}\leq\frac{1}{n}.$
The proof is complete. ∎
2.3 Quasisimple groups
@@
The next two lemmas are essentially well known. We write down the proof of Lemma 2.10 below, which is a little more general, and will be useful later.
@@
Let $Z=\mathbf{Z}(G)$. We first show that $\mathbf{Z}(G/N)=Z/N$. It is clear that $Z/N\subseteq\mathbf{Z}(G/N)$. On the other hand, $\mathbf{Z}(G/N)$ is certainly contained in the projection of $\mathbf{Z}(G/Z)$ in $G/N$. By Grün’s Lemma [12] we have $\mathbf{Z}(G/Z)=1$, so $\mathbf{Z}(G/N)\subseteq Z/N$. Now from the hypotheses we have $N_{0}/N\subseteq Z/N$, i.e. $N_{0}\subseteq Z$. ∎
We conclude this subsection with the following observation, which is useful in an inductive argument.
@@
$h,g]N_{i}=[hN_{i},gN_{i}]\subseteq N_{i}.$
It follows that $h,g]\in N_{1}\cap N_{2}=1$, and since $h\in H$ is arbitrary, we conclude $H\subseteq\mathbf{Z}(G)$. Finally, by Lemma 2.9, $G$ itself is quasisimple, against our initial assumption. ∎
2.4 Linear and projective representations
@@
For the proof of Theorem B, we need detailed information about the $\frac{1}{4}$-quasirandom finite simple groups. The infinite families of $\frac{1}{4}$-quasirandom simple groups are
@@
$G_{2}(q)\hbox{ with }q=3\>(\mathrm{mod}\,6),\quad^{2}B_{2}(q^{2}),\quad^{2}G_{2}(q^{2})\,.$
The relevant information can be recovered from [15, 24] for classical groups, and from [17, Sec. 4] for exceptional groups (see also the Appendix). We collect these data in Table 1.
@@
$\dfrac{1^{1}}{1_{1_{1}}}$ $\mathrm{PSL}_{2}(q)$ $q$ even q1q-1 q1q-1 q1q-1 22
q=1(mod 4)q=1\,($\mathrm${mod}\,4) $\dfrac{q+1}{2}$ $\dfrac{q+1}{2}$ $\dfrac{q-1}{2}$ 22
q=3(mod 4)q=3\,($\mathrm${mod}\,4) $\dfrac{q-1}{2}$ $\dfrac{q-1}{2}$ $\dfrac{q-1}{2}$ 22
$\dfrac{1^{1}}{1_{1_{1}}}$ $\mathrm{PSL}_{3}(q)$ $q\neq 2,4$ $q(q+1)$ $q(q+1)$ $q^{2}+q-1$ 33
$\dfrac{1^{1}}{1_{1_{1}}}$ $\mathrm{Sp}_{4}(q)$ $q$ even $\dfrac{1}{2}q(q-1)^{2}$ $\dfrac{1}{2}q(q-1)^{2}$ $\dfrac{1}{2}q(q-1)^{2}$ 44
$\dfrac{1^{1}}{1_{1_{1}}}$ $G_{2}(q)$ q=3(mod 6)q=3($\mathrm${mod}\,6) $\phi_{3}(q)\phi_{6}(q)$ $\phi_{3}(q)\phi_{6}(q)$ $\phi_{3}(q)\phi_{6}(q)-\delta_{1}$ 77
$\dfrac{1^{1}}{1_{1_{1}}}$ $^{2}B_{2}(q^{2})$ $q=\sqrt{2^{2m+1}}$ $\dfrac{\sqrt{2}}{2}q\phi_{1}(q)\phi_{2}(q)$ $\dfrac{\sqrt{2}}{2}q\phi_{1}(q)\phi_{2}(q)$ $\dfrac{\sqrt{2}}{2}q\phi_{1}(q)\phi_{2}(q)$ 44
$\dfrac{1^{1}}{1_{1_{1}}}$ $^{2}G_{2}(q^{2})$ $q=\sqrt{3^{2m+1}}$ $\phi_{12}(q)$ $\phi_{12}(q)$ $\phi_{12}(q)-\delta_{2}$ 77
Table 1: The $\frac{1}{4}$-quasirandom simple groups of Lie type.
We denote by $\phi_{\ell}(q)$ the $\ell$-th cyclotomic polynomial in the variable $q$. The value of $\delta_{1}$ is $1$ if $r=2$, $q=3^{f}$, $f\geq 2$, or $r=3$, 33 divides q1q-1, and it is 0 otherwise. The value of $\delta_{2}$ is $1$ if $r=2$, and 0 otherwise.
@@
3 Quasirandom groups of affine type
In this short section, we provide some highly quasirandom nonquasisimple groups. The key tool is Lemma 3.2 below. Before giving the results, we revise some notions from the representation theory of finite groups. Let $V$ be a vector space over $\mathbb{C}$, and let $V^{*}$ be its dual space. We remark that $V^{*}\cong\mathrm{Hom}(V,\mathbb{C}^{*})$. If $\phi\colon V\rightarrow V$ is a homomorphism of vector spaces, then $\phi^{*}\colon V^{*}\rightarrow V^{*}$ is defined by
$\phi^{*}(f)\>:=\>f\circ\phi.$
Let $K$ be a finite group. If $\rho\colon K\rightarrow\mathrm{GL}(V)$ is a representation of $K$, then the dual representation $\rho^{*}\colon K\rightarrow\mathrm{GL}(V^{*})$ is defined by
@@ 3 Quasirandom groups of affine type
4 Proof of the main theorems
4.1 Proof of Theorem A
@@
Fix $\varepsilon,\delta>0$, and let $G$ be a perfect group. Suppose that $G/\mathrm{Sol}(G)$ is $\varepsilon$-quasirandom, and that $|G/\mathrm{Sol}(G)|\geq|G|^{\delta}$. If $|G|$ is sufficiently large with respect to $\varepsilon$ and $\delta$, then $G$ is $\frac{\varepsilon\delta}{3}$-quasirandom.
@@
$\mathbf{D}_{0}(G)\>\geq\>c\,\mathbf{D}_{0}(G/\mathrm{Sol}(G))^{1/2},$
where $c$ is an absolute constant. It follows that
$|G|^{\bm{\varepsilon}(G)}\>\geq\>c|G/\mathrm{Sol}(G)|^{\bm{\varepsilon}(G/\mathrm{Sol}(G))/2},$
@@
$\bm{\varepsilon}(G)\>\geq\>\frac{\bm{\varepsilon}(G/\mathrm{Sol}(G))}{2}\cdot\frac{\log|G/\mathrm{Sol}(G)|}{\log|G|}-\frac{\log(c^{-1})}{\log|G|}.$
@@
The main idea behind the other direction of Theorem A is the following. Under the hypothesis that $\mathrm{Sol}(G)=1$, we prove that $G=F^{*}(G)$, the generalized Fitting subgroup of $G$. If this is not the case, then $G/F^{*}(G)$ has a nontrivial simple quotient $L$. We use the quasirandomness of $L$ to obtain an upper bound on $|G|$, while Bender’s theorem and other considerations provide a lower bound on $|F^{*}(G)|$. These inequalities fight each other, giving the desired contradiction. A similar argument will be used in a portion of the proof of Theorem B. Going into details, we start recording the Schreier conjecture, which is now a theorem as a consequence of the Classification of the Finite Simple Groups.
@@
Let $N\lhd G$, and suppose that a section $H_{0}/N_{0}\cong L$ of $G$ is simple. Then $N\subseteq N_{0}$, or $NN_{0}\geqslant H_{0}$. In the first case $L$ appears as a section of $G/N$, while in the second case $L$ appears as a section of $N$.
@@
After Lemma 4.1, we have only to prove the first part of the theorem. Fix $\varepsilon>0$. From Corollary 2.3(i), it is clear that $|G/\mathrm{Sol}(G)|\geq|G|^{\varepsilon}$. Moreover, $G/\mathrm{Sol}(G)$ is again $\varepsilon$-quasirandom. Up to replacing $G$ with $G/\mathrm{Sol}(G)$, we can assume that $G$ has a trivial solvable radical. Now it is well known that the generalized Fitting subgroup $F=F^{*}(G)$ coincides with the socle of $G$, and it is the direct product of nonabelian simple groups $T_{i}$’s. Let $F=\prod_{i}(T_{i})^{n_{i}}$, where $n_{i}\geq 1$ for each $i$. Moreover, as $\mathbf{C}_{G}(F)\subseteq F$ [1, (31.13)], we have $\mathbf{C}_{G}(F)=\mathbf{Z}(F)=1$. It follows that $G$ can be embedded as a subgroup of
@@
$|F|\>\geq\>60^{\mathbf{D}_{0}(L)}.$
When $\mathbf{D}_{0}(L)$ is sufficiently large, this gives a contradiction to (2). We have proven that $G/\mathrm{Sol}(G)=F^{*}(G/\mathrm{Sol}(G))$, so $G/\mathrm{Sol}(G)$ is an $\varepsilon$-quasirandom direct product of finite simple groups, say $G/\mathrm{Sol}(G)=\prod_{i}(T_{i})^{n_{i}}$. From Remark 4.5, each $T_{i}$ is a simple group of Lie type of $\varepsilon$-bounded rank. Let $N=\sum_{i}n_{i}$. From Lemma 2.8(ii), we have $\varepsilon\leq\bm{\varepsilon}(G/\mathrm{Sol}(G))\leq N^{-1}$, so that $N\leq\varepsilon^{-1}$. ∎
@@
An attentive reader will have noticed that a bound of the type $\mathbf{D}_{0}(G)\geq c\log|G|$ is strong enough to give that $G/\mathrm{Sol}(G)$ is the direct product of nonabelian finite simple groups.
It is natural to ask whether the condition $|G/\mathrm{Sol}(G)|\geq|G|^{\varepsilon}$ is actually superflous in the second part of Theorem A, as it may be implied by the quasirandomness of the sole $G/\mathrm{Sol}(G)$. The answer is no, as it can even happen that $|G/\mathrm{Sol}(G)|$ is bounded: it is enough to consider the perfect deleted permutation module $(C_{n})^{4}\rtimes\mathrm{Alt}(5)$ when $n$ is large. The following example, where $|G/\mathrm{Sol}(G)|$ is unbounded but too small, is due to S. Eberhard.
@@
It can be proven that $G_{d}$ is perfect, and its solvable radical is the subgroup where $g\in\mathbf{Z}(\mathrm{SL}_{2}(p))$ is a scalar matrix. In particular, $G_{d}/\mathrm{Sol}(G_{d})$ is isomorphic to $\mathrm{PSL}_{2}(p)$, which is $0.34$-quasirandom for large $p$, but its cardinality is less than $p^{3}$.
4.2 Proof of Theorem B
We first prove the shallow version of Theorem B. This is a nice application of (1).